# Gradient-Based Multi-Objective Deep Learning: Algorithms, Theories, Applications, and Beyond

### v2

@@ 4.2 Methods without Using Preference Vectors
As HV evaluates solution sets based on both convergence and diversity, it is commonly used in traditional evolutionary MOO [86]. In the context of gradient-based MOO, gradient-based hypervolume maximization algorithms (GradHV) [47, 44, 45, 46] have been developed. Let $\{{\bm{\theta}}^{(1)},\dots,{\bm{\theta}}^{(n)}\}$ be a set of solutions and $\{{\bm{f}}({\bm{\theta}}^{(1)}),\dots,{\bm{f}}({\bm{\theta}}^{(n)})\}$ be their corresponding objective vectors. These algorithm first calculate the hypervolume gradient of the $i$-th solution ${\bm{\theta}}^{(i)}$ as follows:
${\bm{d}}^{(i)}=\sum_{k=1}^{m}\underbrace{\frac{\partial\mathrm{HV}_{\bm{r}}(\{{\bm{f}}({\bm{\theta}}^{(1)}),\dots,{\bm{f}}({\bm{\theta}}^{(n)})\})}{\partial f_{k}({\bm{\theta}}^{(i)})}}_{a_{i}:1\times 1}\underbrace{\frac{\partial f_{k}({\bm{\theta}}^{(i)})}{\partial{\bm{\theta}}^{(i)}}}_{{\bm{B}}:1\times d}.$ (32)
@@
Once ${\bm{d}}^{(i)}$ is calculated, solutions ${\bm{\theta}}^{(i)}$’s are updated by gradient ascent, ${\bm{\theta}}^{(i)}\leftarrow{\bm{\theta}}^{(i)}+\eta{\bm{d}}^{(i)}$, where $\eta$ is a learning rate. The term ${\bm{B}}$ in Equation (32) can be easily estimated via backpropagation. The main challenge lies in estimating the first term $a_{i}$. Following Emmerich et al. [46], index sets are categorized based on differentiability: ${\mathbb{Z}}$ includes subvectors with zero partial gradients, ${\mathbb{U}}$ contains subvectors with undefined or indeterminate gradients, and ${\mathbb{P}}$ has subvectors with positive gradients. This classification leads to many zero-gradient terms, improving efficiency. Then, non-zero terms are computed using a fast dimension-sweeping algorithm [87, 88]. This method applies to bi-, tri-, and four-objective problems with time complexities of $\Theta(mn+n\log n)$ for $m=2$ or $3$ and $\Theta(mn+n^{2})$ for $m=4$, where $m$ is the number of objectives and $n$ is the number of solutions. To further enhance convergence, the Newton-hypervolume method [89] incorporates second-order information for faster updates. Deist et al.  [44] apply the hypervolume gradient methods for deep learning tasks such as medical image classification with more than thousands of parameters.
@@
Inspired by Stein Variational Gradient Descent (SVGD) [90], MOO-SVGD [48] proposes a different approach to maintain diversity while enhancing convergence. The update direction for the $i$-th solution ${\bm{\theta}}^{(i)}$ is given by:
${\bm{d}}^{(i)}=$ $-\sum_{j=1}^{n}{\bm{g}}^{*}({\bm{\theta}}^{(j)})\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))$
$\qquad\qquad+\mu\nabla_{{\bm{\theta}}^{(i)}}\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)})).$ (33)
@@
where ${\bm{g}}^{*}({\bm{\theta}}^{(j)})=\sum_{i=1}^{m}\lambda_{i}^{*}\nabla f_{i}({\bm{\theta}}^{(j)}),$ and $\{\lambda_{i}^{*}\}_{i=1}^{m}$ are the objective weights obtained by MGDA [6] (i.e., problem (13)). $\kappa(\cdot,\cdot)$ represents a kernel function, often chosen as the radial basis function (RBF) kernel. The first term in (33) guides the solution towards the PS, while the second term pushes the solutions away from each other to enhance the diversity.
@@
The methods discussed above focus on scenarios where the number of solutions $n$ exceeds the number of objectives $m$. Recent works [49, 51, 91, 50] explore the case with more objectives than solutions (i.e., $m>n$). These approaches assign each objective to a solution ${\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}$ that achieves the smallest objective value for that particular objective. They then minimize either the maximum or the sum of these smallest objective values. Few for Many (F4M) [49, 91] uses the maximum aggregation, formulated as:
$\min_{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}}\max_{i\in[m]}\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}}).$ (34)
Sum of Minimum (SoM) [50] uses the sum aggregation, formulated as:
$\min_{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}}\sum_{i=1}^{m}\left(\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}})\right).$ (35)
@@
MosT [51] frames the problem as a bilevel optimization task, with the outer loop using MGDA to optimize solutions and the inner loop solving an optimal transport problem to assign objectives to solutions.
4.3 Discussions
This section reviews two primary approaches for obtaining multiple Pareto optimal solutions.
@@ 4.3 Discussions
Preference vector-based methods decompose an MOO problem into multiple independent subproblems based on predefined preference vectors. These subproblems can be solved independently, which enhances memory efficiency and enables parallel computation. However, in real-world applications, determining suitable preference vectors can be challenging due to the complex and often unpredictable nature of the Pareto front. Furthermore, since the subproblems are solved independently, they cannot benefit from shared information between one another.
@@
5 Finding an Infinite Set of Solutions
While a finite set of solutions can only provide a discrete approximation of the Pareto front, many applications require the ability to obtain solutions corresponding to any user preference, effectively representing the entire Pareto set. Directly learning an infinite number of solutions individually is impractical.
Ma et al. [92] initially investigated deriving an infinite set of solutions by approximating the Pareto set with first-order expansion around discrete Pareto-optimal solutions. However, this method has several drawbacks: (1) The approximation error grows when the solutions are widely spaced, (2) First-order approximations perform poorly in high-dimensional objective spaces, and (3) The approach requires solving a linear programming problem for updates, which can be computationally challenging.
To address these issues, many methods have been introduced that leverage neural networks to learn mappings from user preferences to solutions directly, enabling the capture of the entire PS. These methods rely on designing efficient network architectures and implementing effective training strategies.
@@ 5 Finding an Infinite Set of Solutions
5.1 Network Structure
@@
Figure 5: Illustration of different structures to learn an infinite number of solutions.
5.1.1 Hypernetwork
Hypernetwork is a neural network that generates the parameters for another target network [93]. This idea has been leveraged by the Pareto Hypernetwork (PHN) [52] and Controllable Pareto Multi-Task Learning (CPMTL) [53] to learn the entire Pareto set, where the input is the user preference ${\bm{\alpha}}$ and the output is the target network’s parameters. The hypernetwork usually consists of several MLP layers, and has been adopted in many subsequent approaches such as PHN-HVI [54] and SUHNPF [55]. Recently, Tuan et al. [56] propose using a transformer architecture as the hypernetwork, which demonstrates superior performance compared to MLP-based hypernetwork models.
A primary limitation of hypernetwork-based algorithms is their sizes. Since the output dimension matches the number of parameters in the target network, hypernetworks are often much larger than the base networks, limiting their applicability to large models. Some methods address this limitation by employing hypernetworks with chunking [52, 53]. Chunking involves partitioning the parameter space into smaller, more manageable segments, enabling the hypernetwork to generate parameters more efficiently and scalability. This approach reduces the overall size of the hypernetwork while preserving its capability to produce accurate parameters for the target network.
5.1.2 Preference-Conditioned Network
Instead of using a hypernetwork to generate weights, the original model can be directly modified to incorporate preferences. COSMOS [57] suggests adding preferences as an additional input to the model by combining user preference ${\bm{\alpha}}$ with input data ${\bm{x}}$, thereby increasing the input dimension of the original neural network.
However, such input-based conditioning has limited ability to produce diverse solutions. Some studies [58, 42, 59] employ the Feature-wise Linear Modulation (FiLM) layer [94] to condition the network. The FiLM layer works by applying an affine transformation to the feature maps. Specifically, given a feature map ${\bm{u}}$ with $C$ channels, we use an MLP to generate conditioning parameters ${\bm{\gamma}}\in\mathbb{R}^{C}$ and ${\bm{\beta}}\in\mathbb{R}^{C}$ based on the given preference ${\bm{\alpha}}$. Then, the FiLM layer modifies the features as ${\bm{u}}^{\prime}_{c}=\gamma_{c}\cdot{\bm{u}}_{c}+\beta_{c},$ where ${\bm{u}}_{c}$ is the $c$-th channel of ${\bm{u}}$. This method achieves better conditioning ability compared to merely adding the preference to the input layer. Raychaudhuri et al. [59] also propose leveraging another network to predict the network architecture. It dynamically adapts the model’s structure according to user preferences, allowing for more flexible and efficient learning.
5.1.3 Model Combination
Methods based on model combination construct a composite model by integrating multiple individual models, thereby offering an effective way to introduce diversity. PaMaL [60] achieves this by learning several base models and combining them through a weighted sum of their parameters based on user preferences:
${\bm{\theta}}({\bm{\alpha}})=\sum_{i=1}^{m}\alpha_{i}{\bm{\theta}}_{i}.$ (36)
@@ 5.1.3 Model Combination
Despite its benefits, learning multiple models simultaneously may become inefficient when handling a large number of objectives. To mitigate this issue, LORPMAN [61] proposes learning a full-rank model ${\bm{\theta}}_{0}\in\mathbb{R}^{p\times q}$ and $m$ low-rank models ${\bm{B}}_{i}{\bm{A}}_{i}$’s, where ${\bm{B}}_{i}\in\mathbb{R}^{p\times r}$ and ${\bm{A}}_{i}\in\mathbb{R}^{r\times q}$. Given user preference ${\bm{\alpha}}$, the composite model is given by:
${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+s\sum_{i=1}^{m}\alpha_{i}{\bm{B}}_{i}{\bm{A}}_{i},$ (37)
where $s$ is a scaling factor that adjusts the impact of the low-rank components. This approach efficiently scales Pareto set learning to manage up to $40$ objectives, achieving superior performance. A similar approach is also mentioned in [65]. To further enhance parameter efficiency, Chen and Kwok [62] introduce a preference-aware tensor multiplication:
${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+s{\bm{C}}\times_{1}{\bm{A}}\times_{2}{\bm{B}}\times_{3}{\bm{\alpha}},$ (38)
where ${\bm{C}}\in\mathbb{R}^{r\times r\times m}$ is a learnable core tensor, ${\bm{A}}\in\mathbb{R}^{r\times p}$, ${\bm{B}}\in\mathbb{R}^{r\times q}$ are learnable matrices, and ×$\times_{u}$ denotes mode-$u$ tensor multiplication [95]. Zhong et al.[63] address scalability challenges for large-scale models, such as language models, using SVD:
${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+{\bm{U}}{\bm{\Sigma}}{\bm{V}},$ (39)
where ${\bm{U}}\in\mathbb{R}^{p\times(r+m)}$ and ${\bm{V}}\in\mathbb{R}^{(r+m)\times q}$ are learnable matrices. ${\bm{\Sigma}}$ is a diagonal matrix defined as $\text{diag}(\sigma_{1},\dots,\sigma_{r},s\alpha_{1},\dots,s\alpha_{m})$, where $\{\sigma_{i}\}_{i=1}^{r}$ and $s$ are learnable scalars.
Instead of training new models, Tang et al. [64] propose to utilize a mixture of experts [96] to combine pre-existing models in order to achieve the Pareto set. Given $n$ models, ${\bm{\theta}}_{1},\dots,{\bm{\theta}}_{n}$, each trained on different datasets, they are first integrated to a unified base model ${\bm{\theta}}_{0}$. For some modules in the network, differences between the individual models and the base model (i.e., ${\bm{\theta}}_{i}-{\bm{\theta}}_{0}$’s) are maintained. These differences are treated as experts within the framework. Then, a mixture of experts is applied, where a gating network conditioned on user preferences determines the weighting of these experts. It enables the generation of diverse models by applying different weighted combinations of experts to align with specific user preferences.
5.2 Training Strategy

### v3

@@ 4.2. Methods without Using Preference Vectors
Unlike methods that rely on preference vectors, another approach directly optimizes for a diverse set of Pareto-optimal solutions. As HV evaluates solution sets based on both convergence and diversity, it is commonly used in traditional evolutionary MOO (Shang et al., 2020). In the context of gradient-based MOO, gradient-based hypervolume maximization algorithms (GradHV) (Wang et al., 2017; Deist et al., 2021, 2020; Emmerich et al., 2007) have been developed. Let $\{{\bm{\theta}}^{(1)},\dots,{\bm{\theta}}^{(n)}\}$ be a set of solutions and $\{{\bm{f}}({\bm{\theta}}^{(1)}),\dots,{\bm{f}}({\bm{\theta}}^{(n)})\}$ be their corresponding objective vectors. These algorithm first calculate the hypervolume gradient of the $i$-th solution ${\bm{\theta}}^{(i)}$ as follows:
(32) ${\bm{d}}^{(i)}=\sum_{k=1}^{m}\underbrace{\frac{\partial\mathrm{HV}_{\bm{r}}(\{{\bm{f}}({\bm{\theta}}^{(1)}),\dots,{\bm{f}}({\bm{\theta}}^{(n)})\})}{\partial f_{k}({\bm{\theta}}^{(i)})}}_{a_{i}:1\times 1}\underbrace{\frac{\partial f_{k}({\bm{\theta}}^{(i)})}{\partial{\bm{\theta}}^{(i)}}}_{{\bm{B}}:1\times d}.$
@@
Once ${\bm{d}}^{(i)}$ is calculated, solutions ${\bm{\theta}}^{(i)}$’s are updated by gradient ascent, ${\bm{\theta}}^{(i)}\leftarrow{\bm{\theta}}^{(i)}+\eta{\bm{d}}^{(i)}$, where $\eta$ is a learning rate. The term $\bm{B}$} in Equation (32) can be easily estimated via backpropagation. The main challenge lies in estimating the first term $a_{i}$. As proposed by Emmerich et al. (Emmerich et al., 2007), index sets are categorized based on differentiability: $\mathbb{Z}$} includes subvectors with zero partial gradients, $\mathbb{U}$} contains subvectors with undefined or indeterminate gradients, and $\mathbb{P}$} has subvectors with positive gradients. This classification leads to many zero-gradient terms, improving efficiency. Then, non-zero terms are computed using a fast dimension-sweeping algorithm (Beume et al., 2009; Guerreiro et al., 2012). This method applies to bi-, tri-, and four-objective problems with time complexities of $\Theta(mn+n\log n)$ for $m=2$ or $3$ and $\Theta(mn+n^{2})$ for $m=4$, where $m$ is the number of objectives and $n$ is the number of solutions. To further enhance convergence, the Newton-hypervolume method (Sosa Hernández et al., 2020) incorporates second-order information for faster updates. Deist et al.  (Deist et al., 2021) apply the hypervolume gradient methods for deep learning tasks such as medical image classification with more than thousands of parameters.
@@
Inspired by Stein Variational Gradient Descent (SVGD) (Liu and Wang, 2016), MOO-SVGD (Liu et al., 2021d) proposes a different approach to maintain diversity while enhancing convergence. The update direction for the $i$-th solution ${\bm{\theta}}^{(i)}$ is given by:
${\bm{d}}^{(i)}=$ $-\sum_{j=1}^{n}{\bm{g}}^{*}({\bm{\theta}}^{(j)})\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))+\mu\nabla_{{\bm{\theta}}^{(i)}}\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)})).$
@@
where ${\bm{g}}^{*}({\bm{\theta}}^{(j)})=\sum_{i=1}^{m}\lambda_{i}^{*}\nabla f_{i}({\bm{\theta}}^{(j)}),$ and $\{\lambda_{i}^{*}\}_{i=1}^{m}$ are the objective weights obtained by MGDA (Sener and Koltun, 2018) (i.e., problem (13)). $\kappa(\cdot,\cdot)$ represents a kernel function, often chosen as the radial basis function (RBF) kernel. The first term in Equation (4.2) drives the particles toward the target distribution, while the second term acts as a repulsive force that promotes diversity.
@@
While the methods discussed above address scenarios with more solutions than objectives ($n>m$), recent work has begun to explore the converse case, where objectives outnumber solutions ($m>n$) (Lin et al., 2025c; Li et al., 2025; Liu et al., 2024b; Ding et al., 2024). These approaches assign each objective to a solution ${\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}$ that achieves the smallest objective value for that particular objective. The goal is then to minimize an aggregation of these smallest objective values, typically either their sum or their maximum. The Few for Many (F4M) framework (Lin et al., 2025c; Liu et al., 2024b), for example, employs a minimax strategy by minimizing the maximum value, which is formulated as:
(33) $\min_{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}}\max_{i\in[m]}\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}}).$
Different from F4M, Sum of Minimum (SoM) (Ding et al., 2024) uses the sum aggregation, formulated as:
(34) $\min_{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}}\sum_{i=1}^{m}\left(\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}})\right).$
@@
MosT (Li et al., 2025) frames the problem as a bilevel optimization task, with the outer loop using MGDA to optimize solutions and the inner loop solving an optimal transport problem to assign objectives to solutions.
4.3. Discussions
@@
This section discusses two primary approaches for obtaining multiple Pareto optimal solutions. Preference vector-based methods decompose a MOP into multiple independent subproblems based on predefined preference vectors. These subproblems can be solved independently, which enhances memory efficiency and enables parallel computation. However, in real-world applications, determining suitable preference vectors can be challenging due to the complex and often unpredictable nature of the Pareto front. Furthermore, since the subproblems are solved independently, they cannot benefit from shared information between one another.
@@
5. Finding an Infinite Set of Solutions
While a finite set of solutions can only provide a discrete approximation of the Pareto front, many applications require the ability to obtain solutions corresponding to any user preference, effectively representing the entire Pareto set. Directly learning an infinite number of solutions individually is impractical. Ma et al. (Ma et al., 2020) initially investigated deriving an infinite set of solutions by approximating the Pareto set with first-order expansion around discrete Pareto-optimal solutions. However, this method has several drawbacks: (1) The approximation error grows when the solutions are widely spaced, (2) First-order approximations perform poorly in high-dimensional objective spaces, and (3) The approach requires solving a linear programming problem for updates, which can be computationally challenging. To address these issues, many methods have been introduced that leverage neural networks to learn mappings from user preferences to solutions directly, enabling the capture of the entire PS. These methods rely on designing efficient network architectures and implementing effective training strategies.
@@ 5. Finding an Infinite Set of Solutions
5.1. Network Structure
5.1.1. Hypernetwork
Hypernetwork is a neural network that generates the parameters for another target network (Ha et al., 2016). This idea has been leveraged by the Pareto Hypernetwork (PHN) (Navon et al., 2021) and Controllable Pareto Multi-Task Learning (CPMTL) (Lin et al., 2020) to learn the entire Pareto set, where the input is the user preference $\bm{\alpha}$} and the output is the target network’s parameters. The hypernetwork usually consists of several MLP layers, and has been adopted in many subsequent approaches such as PHN-HVI (Hoang et al., 2023) and SUHNPF (Gupta et al., 2022). Recently, Tuan et al. (Tuan et al., 2024) propose using a transformer architecture as the hypernetwork, demonstrating superior performance compared to MLP-based hypernetwork models.
A primary limitation of hypernetwork-based algorithms is their sizes. Since the output dimension matches the number of parameters in the target network, hypernetworks are often much larger than the base networks, limiting their applicability to large models. Some methods address this limitation by employing hypernetworks with chunking (Navon et al., 2021; Lin et al., 2020). Chunking involves partitioning the parameter space into smaller, more manageable segments, enabling the hypernetwork to generate parameters more efficiently and scalability. This approach reduces the overall size of the hypernetwork while preserving its capability to produce accurate parameters for the target network.
5.1.2. Preference-Conditioned Network
Instead of using a hypernetwork to generate weights, the original model can be directly modified to incorporate preferences. COSMOS (Ruchte and Grabocka, 2021) suggests adding preferences as an additional input to the model by combining user preference $\bm{\alpha}$} with input data $\bm{x}$}, thereby increasing the input dimension of the original model.
However, such input-based conditioning has limited ability to produce diverse solutions. Some studies (Dosovitskiy and Djolonga, 2020; Chen and Kwok, 2022; Raychaudhuri et al., 2022) employ the Feature-wise Linear Modulation (FiLM) layer (Perez et al., 2018) to condition the network. The FiLM layer works by applying an affine transformation to the feature maps. Specifically, given a feature map $\bm{u}$} with $C$ channels, we use an MLP to generate conditioning parameters ${\bm{\gamma}}\in\mathbb{R}^{C}$ and ${\bm{\beta}}\in\mathbb{R}^{C}$ based on the given preference $\bm{\alpha}$}. Then, the FiLM layer modifies the features as ${\bm{u}}^{\prime}_{c}=\gamma_{c}\cdot{\bm{u}}_{c}+\beta_{c},$ where ${\bm{u}}_{c}$ is the $c$-th channel of $\bm{u}$}. This method achieves better conditioning ability compared to merely adding the preference to the input layer. Raychaudhuri et al. (Raychaudhuri et al., 2022) also propose leveraging another network to predict the network architecture. It dynamically adapts the model’s structure according to user preferences, allowing for more flexible and efficient learning.
@@ 5.1.2. Preference-Conditioned Network
Figure 6. Illustration of different structures to learn an infinite number of solutions.
5.1.3. Model Combination
Methods based on model combination construct a composite model by integrating multiple individual models, thereby offering an effective way to introduce diversity. PaMaL (Dimitriadis et al., 2023) achieves this by learning several base models and combining them through a weighted sum of their parameters based on user preferences:
(35) ${\bm{\theta}}({\bm{\alpha}})=\sum_{i=1}^{m}\alpha_{i}{\bm{\theta}}_{i}.$
@@ 5.1.3. Model Combination
Despite its benefits, learning multiple models simultaneously may become inefficient when handling a large number of objectives. To mitigate this issue, LORPMAN (Chen and Kwok, 2024a) proposes learning a full-rank model ${\bm{\theta}}_{0}\in\mathbb{R}^{p\times q}$ and $m$ low-rank models ${\bm{B}}_{i}{\bm{A}}_{i}$’s, where ${\bm{B}}_{i}\in\mathbb{R}^{p\times r}$ and ${\bm{A}}_{i}\in\mathbb{R}^{r\times q}$. Given user preference $\bm{\alpha}$}, the composite model is given by:
(36) ${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+s\sum_{i=1}^{m}\alpha_{i}{\bm{B}}_{i}{\bm{A}}_{i},$
where $s$ is a scaling factor that adjusts the impact of the low-rank components. This approach efficiently scales Pareto set learning to manage up to $40$ objectives, achieving superior performance. A similar approach is also mentioned in (Dimitriadis et al., 2025). To further enhance parameter efficiency, Chen and Kwok (Chen and Kwok, 2024b) introduce a preference-aware tensor multiplication:
(37) ${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+s{\bm{C}}\times_{1}{\bm{A}}\times_{2}{\bm{B}}\times_{3}{\bm{\alpha}},$
where ${\bm{C}}\in\mathbb{R}^{r\times r\times m}$ is a learnable core tensor, ${\bm{A}}\in\mathbb{R}^{r\times p}$, ${\bm{B}}\in\mathbb{R}^{r\times q}$ are learnable matrices, and $\times_{u}$ denotes mode-$u$ tensor multiplication (Kolda and Bader, 2009). Zhong et al. (Zhong et al., 2024) address scalability challenges for large-scale models, such as large language models, using SVD-LoRA:
(38) ${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+{\bm{U}}{\bm{\Sigma}}{\bm{V}},$
where $\bm{\Sigma}$} is a diagonal matrix defined as $\text{diag}(\sigma_{1},\dots,\sigma_{r},s\alpha_{1},\dots,s\alpha_{m})$, $\{\sigma_{i}\}_{i=1}^{r}$ and $s$ are learnable scalars. ${\bm{U}}\in\mathbb{R}^{p\times(r+m)}$ and ${\bm{V}}\in\mathbb{R}^{(r+m)\times q}$ are learnable matrices.
Instead of training new models, Tang et al. (Tang et al., 2024) propose to utilize a mixture of experts (Cai et al., 2024) to combine pre-existing models in order to achieve the Pareto set. Given $n$ models, ${\bm{\theta}}_{1},\dots,{\bm{\theta}}_{n}$, each trained on different datasets, they are first integrated to a unified base model ${\bm{\theta}}_{0}$. For some modules in the network, differences between the individual models and the base model (i.e., ${\bm{\theta}}_{i}-{\bm{\theta}}_{0}$’s) are maintained. These differences are treated as experts within the framework. Then, a mixture of experts is applied, where a gating network conditioned on user preferences determines the weighting of these experts. It enables the generation of diverse models by applying different weighted combinations of experts to align with specific user preferences.
5.2. Training Strategy